(1 point) Consider the linear system ; = -5 -3): a. Find the eigenvalues and eigenvectors for the coefficient matrix. 21 = vPLEASE ANSWER AND FILL IN ALL ANSWER BOXES PLEASE ANSWER ALL QUESTIONS ASKED
Answer #1

Now, The ☺ 3 2 -5 -3 [] given System is> } = Ay where } - Now, 1A-21) = 0 = (3-1)(-3-1) + 10 = 0. (2+3) (9-3) +10 = 0. → 22 -*3-(3+0) (3+1) y, 2 Now (3+i) 9, +24 - 0 -54, +(-3+1)%=0 If y=2d. then where d ER-{o} * = -(3+i)d Then, 2 & 22=(-1) -sto] tak(i) — Lij) 26 = (-2+41) G = (-1+20) Therefore, 9, (+) = 29 eit +2gelt ->y, (t) = fag (-41-2) (lost + usint) + (+2+41) (6054 E